Suppose A is a 3x3 matrix such that det(A) =16.Find the value of det[(4(A^-1)^T]


1
Expert's answer
2020-10-07T17:44:52-0400

Taking into account that "\det(M^{-1})=\frac{1}{\det (M)}" and "\det(M^T)=\det (M)" for any matrix "M", we conclude that "\det ((A^{-1})^T)=\det (A^{-1})=\frac{1}{\det A}=\frac{1}{16}."

Since "\det(kM)=k^3\det(M)" for any "3\times 3" matrix "M",

"\det (4(A^{-1})^T)=4^3\det ((A^{-1})^T)=4^3\frac{1}{16}=4."


Answer: "4."